OpenStudy (anonymous):

solutions to x^2+10x+25=8

OpenStudy (anonymous):

u need to use quadratic formula only. x^2+10x+17=0

OpenStudy (anonymous):

(8/x+5)-5

OpenStudy (anonymous):

I still don't understand

OpenStudy (anonymous):

Is one of them -sqrt 17/10

OpenStudy (anonymous):

you can write the LHS as a square: (x+5)^2=8 square root both sides x+5=+/-sqrt(8) subtract 5 x=2sqrt(2)-5,-2sqrt(2)-5